Given the following data, answer the questions below the table: -------------------------------------------------------------------------------------------------------------------------------- Data Deluxe Tourist Annual sales...

  1. Given the following data, answer the questions below the table:

--------------------------------------------------------------------------------------------------------------------------------

Data

Deluxe

Tourist

Annual sales in units

2,000

10,000

Direct materials per unit

$25

$17

Direct labor-hours per unit

5

4

Direct labor rate

$12

per DLH

Estimated

Overhead

Expected Activity

Activities and Activity Measures

Cost

Deluxe

Tourist

Total

Labor related (direct labor-hours)

$   80,000

10,000

40,000

50,000

Machine setups (setups)

150,000

3,000

2,000

5,000

Production orders (orders)

70,000

230

270

500

General factory (machine-hours)

250,000

12,000

28,000

40,000

$ 550,000

--------------------------------------------------------------------------------------------------------------------------------

What is the product cost for the Deluxe model?

What is the Activity Based Costing unit product cost for the Deluxe model? (Round your final answer to 2 decimal places.)
Answer #1
Production Cost (ABC)
Deluxe Tourist
Direct Material given 2000*25 = 50000 10,000*17 = 170000
Direct Labor 2000*5*12 = 120,000 10000*4*12= 480000
Prime Cost 170,000 650,000
Labor related 80,000*1/5 = 16000 80000*4/5 = 64000
Machine set ups 150000*3/5 = 90,000 15000*2/5 = 60000
Production orders 70000*230/500 = 32200 70000*270/500 = 37800
General factory 250000*12/40 = 75000 250000*28/40 = 175000
Total product cost 383200 986800

Traditional Costing:

Total DLH: 2000 units * 5 + 10000 units * 4 = 50,000 DLH

So, Overhead rate per DLH = 550,000/50,000 =:11$

Deluxe Tourist
Prime cost 170000 650000
Overhead 2000*5*11 = 110000 10,000*4*11 = 440000
Total Product Cost 280000 1090000